# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

### v2

@@ 3.2 Basic Network
BEV Encoder. Given the fused feature , we further refine by passing it through three blocks of ResNet18 [12], resulting in two feature maps, and , extracted at two scales from the first and last blocks respectively. Then, similar to FPN [27], we apply bilinear upsampling to and concatenate it with along the feature dimension. Finally, we fuse the features of different scales using a convolution block. The output refined BEV features can be denoted as .
3.3 BVRE
@@
3.4 Detachable Auxiliary Detection Head
@@
3.5 Overall Objective Function
Most existing methods rely on complex combinations of loss functions [49, 50, 24, 58, 42, 53, 37, 61, 38, 43], such as a combination of several or all of the focal loss [41], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [22], etc., to achieve the expected performance. In contrast, our approach utilizes a detachable 3D object detection branch for auxiliary supervision and requires only a simple cross-entropy loss for occupancy prediction. The total loss for our framework can be defined as:
4 Experiments
Method | mIoU | Modality | Image Size | 2D Backbone |
@@ 4 Experiments
BEVDetOcc(2f) [13] | 42.02 | C | 512 1408 | Swin-B | 12.2 | 49.6 | 25.1 | 52.0 | 54.5 | 27.9 | 28.0 | 28.9 | 27.2 | 36.4 | 42.2 | 82.3 | 43.3 | 54.6 | 57.9 | 48.6 | 43.6 |
PanoOcc(4f) [50] | 42.13 | C | 864 1600 | R101-DCN | 11.7 | 50.5 | 29.6 | 49.4 | 55.5 | 23.3 | 33.3 | 30.6 | 31.0 | 34.4 | 42.6 | 83.3 | 44.2 | 54.4 | 56.0 | 45.9 | 40.4 |
FB-OCC(16f) [24] | 48.90 | C | 960 1760 | VoVNet-99 | 14.3 | 57.0 | 38.3 | 57.7 | 62.1 | 34.4 | 39.4 | 38.8 | 39.4 | 42.9 | 50.0 | 86.0 | 50.2 | 60.1 | 62.5 | 52.4 | 45.7 |
FlashOcc(2f) [57] | 43.52 | C | 512 1408 | Swin-B | 13.4 | 51.1 | 27.7 | 51.6 | 56.2 | 27.3 | 30.0 | 29.9 | 29.8 | 37.8 | 43.5 | 83.8 | 46.6 | 56.2 | 59.6 | 50.8 | 44.7 |
COTR(8f) [36] | 46.21 | C | 512 1408 | Swin-B | 14.9 | 53.3 | 35.2 | 50.8 | 57.3 | 35.4 | 34.1 | 33.5 | 37.1 | 39.0 | 45.0 | 84.5 | 48.7 | 57.6 | 61.1 | 51.6 | 46.7 |
RadOcc [60] | 49.38 | C+L | 512 1408 | Swin-B | 10.9 | 58.2 | 25.0 | 57.9 | 62.9 | 34.0 | 33.5 | 50.1 | 32.1 | 48.9 | 52.1 | 82.9 | 42.7 | 55.3 | 58.3 | 68.6 | 66.0 |
OccFusion [37] | 46.67 | C+L+R | 900 1600 | R101-DCN | 12.4 | 50.3 | 31.5 | 57.6 | 58.8 | 34.0 | 41.0 | 47.2 | 29.7 | 42.0 | 48.0 | 78.4 | 35.7 | 47.3 | 52.7 | 63.5 | 63.3 |
OccFusion [61] | 48.74 | C+L | 900 1600 | R101 | 12.4 | 51.8 | 33.0 | 54.6 | 57.7 | 34.0 | 43.0 | 48.4 | 35.5 | 41.2 | 48.6 | 83.0 | 44.7 | 57.1 | 60.0 | 62.5 | 61.3 |
HyDRa [52] | 44.4 | C+R | 256 704 | R50 | - | - | - | 52.3 | 56.3 | - | 35.9 | 35.1 | - | - | 44.1 | - | - | - | - | - | - |
EFFOcc [43] | 52.62 | C+L | 512 1408 | Swin-B | 14.7 | 58.6 | 34.7 | 60.5 | 65.5 | 36.0 | 40.4 | 51.5 | 41.0 | 48.5 | 54.6 | 84.2 | 46.4 | 57.9 | 60.8 | 70.9 | 68.4 |
DAOcc (Ours) | 53.82 | C+L | 256 704 | R50 | 12.4 | 59.6 | 38.4 | 61.9 | 67.1 | 35.3 | 48.2 | 59.1 | 43.5 | 50.9 | 56.3 | 83.0 | 44.7 | 56.7 | 59.9 | 70.0 | 68.1 |
Method | RayIoU | Modality | Image Size | 2D Backbone | RayIoU1m | RayIoU2m | RayIoU4m | mIoU |
BEVFormer(4f) [23] | 32.4 | C | 900 1600 | R101 | 26.1 | 32.9 | 38.0 | 39.2 |
RenderOcc [39] | 19.5 | C | 512 1408 | Swin-B | 13.4 | 19.6 | 25.5 | 24.4 |
SimpleOcc [9] | 22.5 | C | 336 672 | R101 | 17.0 | 22.7 | 27.9 | 31.8 |
BEVDetOcc(8f) [13] | 32.6 | C | 384 704 | R50 | 26.6 | 33.1 | 38.2 | 39.3 |
FB-Occ(16f) [24] | 33.5 | C | 256 704 | R50 | 26.7 | 34.1 | 39.7 | 39.1 |
SparseOcc(16f) [31] | 36.1 | C | 256 704 | R50 | 30.2 | 36.8 | 41.2 | 30.9 |
Panoptic-FlashOcc [58] | 35.2 | C | 256 704 | R50 | 29.4 | 36.0 | 40.1 | 29.4 |
Panoptic-FlashOcc(8f) [58] | 38.5 | C | 256 704 | R50 | 32.8 | 39.3 | 43.4 | 31.6 |
DAOcc (Ours) | 48.2 | C+L | 256 704 | R50 | 44.4 | 48.7 | 51.6 | 47.9 |
Method | IoU | mIoU | Modality | Image Size | 2D Backbone |
@@
SurroundOcc [51] | 31.4 | 20.3 | C | 900 1600 | R101-DCN | 20.5 | 11.6 | 28.1 | 30.8 | 10.7 | 15.1 | 14.0 | 12.0 | 14.3 | 22.2 | 37.2 | 23.7 | 24.4 | 22.7 | 14.8 | 21.8 |
OccFormer [62] | 29.9 | 20.1 | C | 896 1600 | R101 | 21.1 | 11.3 | 28.2 | 30.3 | 10.6 | 15.7 | 14.4 | 11.2 | 14.0 | 22.6 | 37.3 | 22.4 | 24.9 | 23.5 | 15.2 | 21.1 |
C-CONet [49] | 26.1 | 18.4 | C | 896 1600 | R101 | 18.6 | 10.0 | 26.4 | 27.4 | 8.6 | 15.7 | 13.3 | 9.7 | 10.9 | 20.2 | 33.0 | 20.7 | 21.4 | 21.8 | 14.7 | 21.3 |
FB-Occ [24] | 31.5 | 19.6 | C | 896 1600 | R101 | 20.6 | 11.3 | 26.9 | 29.8 | 10.4 | 13.6 | 13.7 | 11.4 | 11.5 | 20.6 | 38.2 | 21.5 | 24.6 | 22.7 | 14.8 | 21.6 |
RenderOcc [39] | 29.2 | 19.0 | C | 896 1600 | R101 | 19.7 | 11.2 | 28.1 | 28.2 | 9.8 | 14.7 | 11.8 | 11.9 | 13.1 | 20.1 | 33.2 | 21.3 | 22.6 | 22.3 | 15.3 | 20.9 |
LMSCNet [40] | 36.6 | 14.9 | L | - | - | 13.1 | 4.5 | 14.7 | 22.1 | 12.6 | 4.2 | 7.2 | 7.1 | 12.2 | 11.5 | 26.3 | 14.3 | 21.1 | 15.2 | 18.5 | 34.2 |
L-CONet [49] | 39.4 | 17.7 | L | - | - | 19.2 | 4.0 | 15.1 | 26.9 | 6.2 | 3.8 | 6.8 | 6.0 | 14.1 | 13.1 | 39.7 | 19.1 | 24.0 | 23.9 | 25.1 | 35.7 |
M-CONet [49] | 39.2 | 24.7 | C+L | 896 1600 | R101 | 24.8 | 13.0 | 31.6 | 34.8 | 14.6 | 18.0 | 20.0 | 14.7 | 20.0 | 26.6 | 39.2 | 22.8 | 26.1 | 26.0 | 26.0 | 37.1 |
Co-Occ [38] | 41.1 | 27.1 | C+L | 896 1600 | R101 | 28.1 | 16.1 | 34.0 | 37.2 | 17.0 | 21.6 | 20.8 | 15.9 | 21.9 | 28.7 | 42.3 | 25.4 | 29.1 | 28.6 | 28.2 | 38.0 |
OccFusion [37] | 44.7 | 27.3 | C+L+R | 900 1600 | R101-DCN | 27.1 | 19.6 | 33.7 | 36.2 | 21.7 | 24.8 | 25.3 | 16.3 | 21.8 | 30.0 | 39.5 | 19.9 | 24.9 | 26.5 | 28.9 | 40.4 |
DAOcc (Ours) | 45.0 | 30.5 | C+L | 256 704 | R50 | 30.8 | 19.5 | 34.0 | 38.8 | 25.0 | 27.7 | 29.9 | 22.5 | 23.2 | 31.6 | 41.0 | 25.9 | 29.4 | 29.9 | 35.2 | 43.5 |
4.1 Datasets and Evaluation Metrics
Occ3D-nuScenes [44] is a large-scale benchmark dataset for 3D occupancy prediction, which contains 700 scenes for training and 150 scenes for validation. Each scene lasts for 20 seconds, with annotations provided at a frequency of 2 Hz. The dataset covers a perception range of [-40m, 40m] in the and directions, and [-1m, 5.4m] along the -axis, all within the ego coordinate system. This range is discretized into voxels, each with a size of [0.4m, 0.4m, 0.4m]. Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class labeled as ‘others’. The dataset includes RGB images captured by six surrounding cameras and LiDAR point clouds, enabling dense voxel-wise annotations for 3D scene understanding. Additionally, the dataset provides visibility masks for both LiDAR and camera modes, which can be utilized for training purposes.
SurroundOcc [51] is an automatically generated occupancy prediction dataset that does not require extra human annotation. It is built upon the nuScenes [3] and generates dense occupancy ground truth by utilizing existing 3D detection and 3D semantic segmentation labels. The dataset uses the nuScenes training set to train the model and employs the nuScenes validation set for evaluation purposes. The occupancy prediction range is set to [-50m, 50m] for the and axes, and [-5m, 3m] for the -axis, all within the LiDAR coordinate system. The final output 3D occupancy grid has a shape of 200×200×16, with each voxel measuring [0.5m, 0.5m, 0.5m].
Evaluation metrics: Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with camera visible mask provided by Occ3D-nuScenes, ensuring a fair comparison with existing approaches. When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [31], which has been recently introduced. For the SurroundOcc dataset, we report semantic segmentation performance using IoU and mIoU. Here, IoU is defined as the intersection over union of occupied voxels, calculated while ignoring their semantic categories, and it is used as an evaluation metric for the scene completion (SC) task.
4.2 Implementation Details
We use ResNet-50 [12], pre-trained on nuImages [3] by MMDetection [5], as our image backbone and adjust the input image resolution to 256×704. In the LiDAR branch, we aggregate LiDAR sweeps from 10 adjacent timestamps as input and initiate training from scratch. Data augmentation techniques, such as scaling, rotation, and flipping, are applied to the input images, and we also perform random flipping in the BEV space. The loss weights are configured as and . We adopt the AdamW optimizer [35] with a cosine annealing learning rate scheduler, incorporating a warmup phase, and set the initial learning rate to 2e-4. Unless conducting ablation experiments, we train our DAOcc using CBGS [65] for 6 epochs, the input point cloud range is set to [-54.0, -54.0, -5.0, 54.0, 54.0, 3.0] and the voxel size is set to [0.075, 0.075, 0.2]. All models are trained on 8 NVIDIA RTX4090 GPUs with a batch size of 4. Additionally, if the camera visible mask is not used during training, we randomly discard 80% of empty voxels, consistent with Zhao et al. [63]. Temporal information fusion, exponential moving average (EMA) for weight parameter and test-time augmentation is not used.
4.3 Comparison with State-of-the-Art Methods
Method | Modality | 3D Backbone | Param | mIoU |
FB-Occ [24] | C | - | 130.8M | 48.90 |
RadOcc [60] | C+L | SECOND | 135.4M | 49.38 |
EFFOcc [43] | C+L | SECOND | 108.6M | 52.62 |
DAOcc | C+L | SECOND | 56.2M | 53.82 |
Comparison on Occ3D-nuScenes with camera mask. As presented in Table 1, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset. All results are provided either directly by their authors or based on their official code implementations. Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 [12] and 256×704 input image resolution. Compared with the latest state-of-the-art multi-modal method EFFOcc [43], DAOcc achieves improvement of 1.2 mIoU. Notably, the image feature extraction network Swin-B [32] in EFFOcc has approximately 3 times the number of parameters of ResNet50, and EFFOcc uses a larger input image resolution of 512×1408. This highlights our method’s ability to mine more useful information from point cloud data. In addition, to achieve optimal results, EFFOcc requires the LiDAR branch to undergo 3D object detection pre-training, which leads to an improvement of approximately 3 mIoU. In contrast, DAOcc does not require LiDAR branch pre-training, but instead seamlessly integrates 3D object detection as an auxiliary branch into the multi-modal occupancy prediction framework, simplifying the model training process and achieving better performance.
Furthermore, we compare the complexity of the models, as shown in Table 4. Compared to the two top-performing multi-modal methods, RadOcc [60] and EFFOcc [43], as well as the top-performing image-based method FB-Occ [24], our proposed DAOcc has only about half the number of parameters but achieves significantly better performance (outperforming RadOcc, EFFOcc, and FB-Occ by 4.44, 1.2 and 4.92 mIoU).
Comparison on Occ3D-nuScenes without camera mask. To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods. As shown in Table 2, benefiting from the effective use of multi-modal input, especially point cloud data, our DAOcc achieves an outstanding 48.2 RayIoU and significantly outperforms image-based methods by a large margin, even though many of these methods use multi-frame temporal fusion. It should be noted that some image-based methods, such as Panoptic-FlashOcc and FB-Occ [24], adopt the Explicit Depth Supervision [22] and leverage depth information derived from point cloud data to supervise the training of monocular depth estimation from images, so these methods can also be regarded as multi-modal during the training phase.
Comparison on SurroundOcc. Table 3 presents a quantitative comparison on the SurroundOcc validation set, showcasing the performance of our DAOcc against other methods. Leveraging both camera and LiDAR inputs, DAOcc achieves a notable 3.2 mIoU improvement over OccFusion [37], which incorporates radar data in addition to camera and LiDAR inputs. It is worth noting that OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower input resolution of 256×704. This underscores the effectiveness and efficiency of our proposed method.
4.4 Ablation Studies
Ablation studies are conducted on the validation set of Occ3D-nuScenes, with the camera visible mask being used during training. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs without employing CBGS [65].
BEV View Range (m) | BEV Voxel Size (m) | mIoU | |
(a) | 0.075 0.075 | 50.76 | |
(b) | 0.075 0.075 | 50.98 | |
(c) | 0.1 0.1 | 49.75 | |
(d) | 0.1 0.1 | 49.98 | |
(e) | 0.075 0.075 | 51.26 |
Effectiveness of the BVRE. We propose the BVRE strategy to enrich spatial contextual information by expanding the perceptual range from a BEV (Bird’s Eye View) perspective. As shown in Table 5, a comparison of (a) and (b) indicates that, with the same voxel size, expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU. Furthermore, the comparison of (a) and (e) shows that further expanding the BEV range to [-54.0m, 54.0m] augments the performance gain to 0.5 mIoU. This observation validates the value of spatial context information in a larger BEV range.
In addition, by comparing (b) and (c), we can find that the size of the voxel has a great impact on the model’s performance. To achieve better performance, we adopt the setting (e) in DAOcc.
camera mask | ADH | mIoU |

### v3

@@ III-B4 BEV Encoder
Given the fused feature , we further refine by passing it through three blocks of ResNet18 [21], resulting in two feature maps, and , extracted at two scales from the first and last blocks respectively. Then, similar to FPN [58], we apply bilinear upsampling to and concatenate it with along the feature dimension. Finally, we fuse the features of different scales using a convolution block. The output refined BEV features can be denoted as .
III-C BVRE
@@
III-D Detachable Auxiliary Detection Head
@@
III-E Overall Objective Function
Most existing methods rely on complex combinations of loss functions [12, 3, 17, 62, 19, 20, 10, 11, 13], such as a combination of several or all of the focal loss [36], the scene-class affinity loss [2], the dice loss, the lovasz-softmax loss [37], the depth loss [38], etc., to achieve the expected performance. In contrast, our approach utilizes a detachable 3D object detection branch for auxiliary supervision and requires only a simple cross-entropy loss for occupancy prediction. The total loss for our framework can be defined as:
IV Experiments
| Method | mIoU | Modality | Image Size | 2D Backbone | FPS |
@@ IV Experiments
| BEVDetOcc(2F) [16] | 42.02 | C | 512 1408 | Swin-B | 1.2 | 12.2 | 49.6 | 25.1 | 52.0 | 54.5 | 27.9 | 28.0 | 28.9 | 27.2 | 36.4 | 42.2 | 82.3 | 43.3 | 54.6 | 57.9 | 48.6 | 43.6 |
| PanoOcc(4F) [3] | 42.13 | C | 864 1600 | R101-DCN | 3.0 | 11.7 | 50.5 | 29.6 | 49.4 | 55.5 | 23.3 | 33.3 | 30.6 | 31.0 | 34.4 | 42.6 | 83.3 | 44.2 | 54.4 | 56.0 | 45.9 | 40.4 |
| FB-OCC(16F) [17] | 48.90 | C | 960 1760 | VoVNet-99 | - | 14.3 | 57.0 | 38.3 | 57.7 | 62.1 | 34.4 | 39.4 | 38.8 | 39.4 | 42.9 | 50.0 | 86.0 | 50.2 | 60.1 | 62.5 | 52.4 | 45.7 |
| FlashOcc(2F)* [18] | 43.52 | C | 512 1408 | Swin-B | 1.2 | 13.4 | 51.1 | 27.7 | 51.6 | 56.2 | 27.3 | 30.0 | 29.9 | 29.8 | 37.8 | 43.5 | 83.8 | 46.6 | 56.2 | 59.6 | 50.8 | 44.7 |
| COTR(8F) [43] | 46.21 | C | 512 1408 | Swin-B | 0.3 | 14.9 | 53.3 | 35.2 | 50.8 | 57.3 | 35.4 | 34.1 | 33.5 | 37.1 | 39.0 | 45.0 | 84.5 | 48.7 | 57.6 | 61.1 | 51.6 | 46.7 |
| STCOcc(16F) [44] | 45.0 | C | 512 1408 | R50 | 1.4 | 15.2 | 52.3 | 32.2 | 50.5 | 56.5 | 31.7 | 33.9 | 33.4 | 33.8 | 38.9 | 44.9 | 83.9 | 47.4 | 57.1 | 60.1 | 50.6 | 42.7 |
| ALOcc(16F) [8] | 50.6 | C | 512 1408 | Swin-B | 1.5 | 17.0 | 58.3 | 39.7 | 56.6 | 63.2 | 33.2 | 41.3 | 40.3 | 40.8 | 43.8 | 51.0 | 87.0 | 52.7 | 62.0 | 65.2 | 57.7 | 50.9 |
| RadOcc [9] | 49.38 | C+L | 512 1408 | Swin-B | - | 10.9 | 58.2 | 25.0 | 57.9 | 62.9 | 34.0 | 33.5 | 50.1 | 32.1 | 48.9 | 52.1 | 82.9 | 42.7 | 55.3 | 58.3 | 68.6 | 66.0 |
| OccFusion [10] | 46.67 | C+L+R | 900 1600 | R101-DCN | 1.0 | 12.4 | 50.3 | 31.5 | 57.6 | 58.8 | 34.0 | 41.0 | 47.2 | 29.7 | 42.0 | 48.0 | 78.4 | 35.7 | 47.3 | 52.7 | 63.5 | 63.3 |
| OccFusion [11] | 48.74 | C+L | 900 1600 | R101 | - | 12.4 | 51.8 | 33.0 | 54.6 | 57.7 | 34.0 | 43.0 | 48.4 | 35.5 | 41.2 | 48.6 | 83.0 | 44.7 | 57.1 | 60.0 | 62.5 | 61.3 |
| SDGOCC [7] | 51.66 | C+L | - | R50 | 7.5 | 13.2 | 57.8 | 24.3 | 60.3 | 64.3 | 36.2 | 39.4 | 52.4 | 35.8 | 50.9 | 53.7 | 84.6 | 47.5 | 58.0 | 61.6 | 70.7 | 67.7 |
| DAOcc (Ours) | 54.33 | C+L | 256 704 | R50 | 7.8 | 13.0 | 60.7 | 39.8 | 64.0 | 66.5 | 36.3 | 49.0 | 60.1 | 44.3 | 50.7 | 55.9 | 82.9 | 44.6 | 56.8 | 60.6 | 70.1 | 68.3 |
| Method | Framework | Precision | Hardware | mIoU | FPS |
| DAOcc | PyTorch | FP32 | RTX4090 | 54.33 | 7.8 |
| DAOcc | TensorRT | FP16 | RTX4090 | 54.20 | 104.9 |
| DAOcc | TensorRT | FP16+INT8 | Orin | 53.70 | 20.0 |
| Method | mIoU | RayIoU | Modality | Image Size | 2D Backbone | FPS | RayIoU1m | RayIoU2m | RayIoU4m | |
| train | val | |||||||||
| BEVFormer(4F) [39] | 39.2 | 32.4 | - | C | 900 1600 | R101 | - | 26.1 | 32.9 | 38.0 |
| RenderOcc [63] | 24.4 | 19.5 | - | C | 512 1408 | Swin-B | - | 13.4 | 19.6 | 25.5 |
| SimpleOcc [64] | 31.8 | 22.5 | - | C | 336 672 | R101 | - | 17.0 | 22.7 | 27.9 |
| BEVDetOcc(8F) [16] | 39.3 | 32.6 | C+L | C | 384 704 | R50 | - | 26.6 | 33.1 | 38.2 |
| FB-Occ(16F) [17] | 39.1 | 33.5 | C+L | C | 256 704 | R50 | - | 26.7 | 34.1 | 39.7 |
| SparseOcc(16F) [42] | 30.9 | 36.1 | - | C | 256 704 | R50 | - | 30.2 | 36.8 | 41.2 |
| Panoptic-FlashOcc [62] | 29.4 | 35.2 | C+L | C | 256 704 | R50 | - | 29.4 | 36.0 | 40.1 |
| Panoptic-FlashOcc(8F) [62] | 31.6 | 38.5 | C+L | C | 256 704 | R50 | - | 32.8 | 39.3 | 43.4 |
| STCOcc(16F) [44] | - | 42.1 | C+L | C | 512 1408 | R50 | 1.4 | 36.9 | 42.8 | 46.7 |
| ALOcc(16F) [8] | 38.0 | 43.7 | C+L | C | 256 704 | R50 | 6.0 | 37.8 | 44.7 | 48.8 |
| DAOcc (Ours) | 48.3 | 48.4 | C+L | C+L | 256 704 | R50 | 7.8 | 44.5 | 48.9 | 51.9 |
| BEV View Range (m) | BEV Voxel Size (m) | H-Range (m) | mIoU | |
| (a) | 0.075 0.075 | [-6, 4] | 50.57 | |
| (b) | 0.075 0.075 | [-4, 2] | 50.27 | |
| (c) | 0.075 0.075 | [-5, 3] | 50.76 | |
| (d) | 0.075 0.075 | [-5, 3] | 50.98 | |
| (e) | 0.1 0.1 | [-5, 3] | 49.75 | |
| (f) | 0.1 0.1 | [-5, 3] | 49.98 | |
| (g) | 0.075 0.075 | [-5, 3] | 51.26 |
ADH | mIoU |
others |
@@
| ✓ | 51.26 | 13.01 | 56.94 | 20.6 | 61.49 | 65.51 | 35.08 | 38.89 | 55.44 | 36.65 | 48.46 | 53.82 | 83.22 | 45.08 | 57.35 | 60.63 | 70.83 | 68.47 | |||||||||||||||||||||||||||||||||||||
| Scheduler | mIoU |
| EP15 | 52.82 |
| EP24 | 53.39 |
| CBGS(EP6) | 53.82 |
| CBGS(EP6)+LR2 | 54.33 |
| Method | mIoU | Modality | Image Size | 2D Backbone |
GO |
vehicle |
bicyclist |
pedestrian |
sign |
traffic light |
pole |
Cons. Cone |